Thankx Pieter!
What are the next two terms of the sequence 32, -48, 72, -108, . . . ?

Found 2 solutions by carl, khwang:
Answer by carl(7)

simple....
32, 72, (add 40), so the next number is 112
AND THEN
-48, -108, (add 60) so the next number is -168
Got it?
Carl

Answer by khwang(438)

32, -48, 72, -108, . . .
is a geometric series with common ratio
r = -48/32 = 72/-48 = -108/ 72 = -3/2
Let a1 = 32, the general term an of the sequence as
an = a1 r^(n-1) for n>= 1.
Hence, the next two terms are a5 = -108 * (-3/2) = 162,(= 32*(-3/2)^4)
and a6 = 162 * (-3/2) = -243 (= 32*(-3/2)^5)
Forget another wrong answer.
Kenny